OpenStudy (anonymous):

Are my steps right so far? What is the 32nd term of the arithmetic sequence where a1 = -34 and a9 = -122? First term: -34 d = -88

hartnn (hartnn):

d is incorrect.

hartnn (hartnn):

d= a2-a1 and not a9-a1

hartnn (hartnn):

to find 'd' use n'th term formula \(a_n=a_1+(n-1)d\) with a1=-34 n=9 a9=-122 find 'd'

OpenStudy (anonymous):

:( where does a2 come from??

hartnn (hartnn):

thats when a2 is given here, to find 'd' use n'th term formula an=a1+(n−1)d with a1=-34 n=9 a9=-122 find 'd'

OpenStudy (anonymous):

wait in the formula you gave, is that 

hartnn (hartnn):

yes, on the left side of =

OpenStudy (anonymous):

ok:) let me figure this out on my paper

OpenStudy (anonymous):

Where do I find the "an" part?

hartnn (hartnn):

for 9th term, n= 9 so, an = a9 =-122

OpenStudy (anonymous):

ohh!!!! okay!!!!

OpenStudy (anonymous):

now finding the common difference..

OpenStudy (anonymous):

-88?

hartnn (hartnn):

again ?

OpenStudy (anonymous):

The common difference: -88

hartnn (hartnn):

an=a1+(n−1)d with a1=-34 n=9 a9=-122 find 'd' -122 = -34 + 8d -88= 8d d=-11 is what i get.

OpenStudy (anonymous):

-122 - -34 = -88

OpenStudy (anonymous):

how did you get 8d?

hartnn (hartnn):

.....(n-1)*d n=9 .....(9-1)*d ......8d

OpenStudy (anonymous):

huh?? i am so confused by that

OpenStudy (anonymous):

how did we get n - 1 in the first place?